OpenStudy (anonymous):

Solve the equation for y: 4x + 3y = -12

OpenStudy (anonymous):

to solve for y subtract 4x on both sides and get 3y = -12 - 4x then divide both sides by 3 y = (-12-4x)/3 simplify y = -4 -4/3x

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

np glad to help